OpenStudy (anonymous):

please someone help me solve this: one of these has n stages of nodes and more than 200 000 leaves. show that n satisfies the inequality n> log (base10) 200 007- log(base10) 7 /log(base10) 2 hence find the least possible value of n. thank u :)

OpenStudy (amistre64):

n > log (base10) 200,007- log(base10) 7 /log(base10) 2; <-2?

OpenStudy (amistre64):

ohhh....i see it now; wierd format...

OpenStudy (amistre64):

log (base10) is normally just written as "log"

OpenStudy (amistre64):

log(200,007) - log(7) / log(2)

OpenStudy (amistre64):

log(200,007/7) ------------- = 14.80 id say 15 log(2)

OpenStudy (amistre64):

7(2^14 - 1) = 114,681

OpenStudy (amistre64):

n is greater than 15 at least :)

OpenStudy (amistre64):

7 * ((2^15) - 1) = 229,369 is good i mean

OpenStudy (amistre64):

i kept punching in 14 instead of 15 lol

OpenStudy (anonymous):

amistre Ure great...do u ever joining IMO?

OpenStudy (amistre64):

i joined when i was a carpenter :)

OpenStudy (anonymous):

haha..i'm sure u get medal on international math. olympiad?

OpenStudy (anonymous):

huh? 7(2^14-1)=114681?? why did you divide it by 10000? :S

OpenStudy (amistre64):

i dint know i did

OpenStudy (amistre64):

7 times [(2^14) - 1] equals 114681 right?

OpenStudy (amistre64):

log(a)-log(b) log(a/b) ---------- = -------- log(c) log(c)

OpenStudy (amistre64):

internation math olympiad eh...... i wonder what that would be like :)

OpenStudy (amistre64):

...its for high school studentlol

OpenStudy (anonymous):

haha may i ask ur email so we can share about math.

OpenStudy (amistre64):

tony031172@gmail.com

OpenStudy (anonymous):

thx mine tianpradana@gmail.com

OpenStudy (anonymous):

thanks :)